OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 6 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 10 inches. What is the volume of water in the vase? Answer 36π in3 54π in3 66π in3 82π in3

OpenStudy (alexwee123):

agh its easier to explain if i could draw but i suck at drawing on the comp -_-

OpenStudy (anonymous):

will you explain how to get the answer please

OpenStudy (anonymous):

im having alot of trouble

OpenStudy (amistre64):

\[V_c-6V_s\]

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

cause i honestly have no idea

OpenStudy (amistre64):

then the answer isnt going to help out to much

OpenStudy (jlastino):

 So compute the volume of vase (which is cylindrical) Volume of cylinder =pi*r^2*h The diameter is 6 inches divide this by 2 to get the tadius r=3 Since the problem says "The vase is filled with water up to a height of 10 inches." we assume that the height of the vase is 10 h=10 use that in the equation pi*3^2*10 this will give 90pi (but this isn't the amount of water yet) Calculate the volume of one marble (which are spherical in shape) Volume of sphere (4/3)*pi*r^3 r=1 (4/3)*pi*1^3 (4/3)pi then multiply it by 6 (there are 6 marbles 6(4/3)*pi 24/3pi 8pi Then subtract this from the volume of the vase 90pi-8pi =82pi